# Sample and Map from a Single Convex Potential: Generation using Conjugate Moment Measures

### v1

@@ Abstract
A common approach to generative modeling is to split model-fitting into two blocks: define first how to sample noise (e.g. Gaussian) and choose next what to do with it (e.g. using a single map or flows). We explore in this work an alternative route that ties sampling and mapping. We find inspiration in moment measures (Cordero-Erausquin and Klartag, 2015), a result that states that for any measure $\rho$ supported on a compact convex set of $\mathbb{R}^{d}$, there exists a unique convex potential $u$ such that $\rho=\nabla u\,\sharp\,e^{-u}$. While this does seem to tie effectively sampling (from log-concave distribution $e^{-u}$) and action (pushing particles through $\nabla u$), we observe on simple examples (e.g., Gaussians or 1D distributions) that this choice is ill-suited for practical tasks. We study an alternative factorization, where $\rho$ is factorized as $\nabla w^{*}\,\sharp\,e^{-w}$, where $w^{*}$ is the convex conjugate of $w$. We call this approach conjugate moment measures, and show far more intuitive results on these examples. Because $\nabla w^{*}$ is the Monge map between the log-concave distribution $e^{-w}$ and $\rho$, we rely on optimal transport solvers to propose an algorithm to recover $w$ from samples of $\rho$, and parameterize $w$ as an input-convex neural network.
@@ 1 Introduction
A decade after the introduction of GAN  (Goodfellow et al., 2014) and VAE (Kingma and Welling, 2014), the field of generative modeling has grown into one of the most important areas of research in machine learning. Both canonical approaches follow the template of learning a transformation that can map random codes to meaningful data. These transformations can be learned in supervised manner, as in the dimensionality reduction pipeline advocated in VAEs, or distributionally as advocated in the purely generative literature. In the latter, the variety of such transforms has gained remarkably in both complexity, using increasingly creative inductive biases, from Neural-ODEs(Chen et al., 2018; Grathwohl et al., 2018), diffusions (Song et al., 2020; Ho et al., 2020), optimal transport (Korotin et al., 2020) to flow-matching (Lipman et al., 2023; Tong et al., 2023; Pooladian et al., 2023).
Sample first, move next. All of these approaches share, however, a standard Gaussian multivariate distribution to generate noise/codes. For both GANs and VAEs, that choice is usually made because of its simplicity, but also, in the case of diffusion models, because Gaussian distributions can be recovered quickly with suitable stochastic processes (Ohrstein-Uhlenbeck). Other works have considered optimizing prior noise distributions (Tomczak and Welling, 2018; Lee et al., 2021; Liang et al., 2022) but still do so in a two step approach where mappings are estimated independently. We propose in this work a new approach where sampling and transforms are not treated separately, but seen as two facets of the same single object.
Tying sampling and action. Our work looks to moment measures Cordero-Erausquin and Klartag (2015) to find inspiration for a factorization that ties both sampling and action.  (Cordero-Erausquin and Klartag, 2015) proved that for any sufficiently regular probability measure $\rho$, one can find an essentially-continuous, convex potential $u$ such that $\rho=\nabla u\,\sharp\,e^{-u}$: essentially, any probability distribution can be recovered by first sampling points from a log-concave distribution $e^{-u}$, and moving them with $\nabla u$ where, remarkably, the same single convex potential $u$ is used twice.
@@
  • After recalling in detail the contribution of  Cordero-Erausquin and Klartag (2015) in §2.1, we argue in §3.1 that the moment measure factorization may not be suitable for practical tasks. For instance, in the case of Gaussian distributions, we show that if $\rho$ is Gaussian with variance $\Sigma$, the corresponding log-concave distribution $e^{-u}$ is Gaussian with variance $\Sigma^{-1}$, amplifying beyond necessary any minor degeneracy in $\Sigma$. Similarly, for univariate $\rho$, we found that the spread of $e^{-u}$ is inversely proportional to that of $\rho$. While these results are not surprising from a theoretical perspective, they suggest a different strategy to estimate a tied sample/action factorization.
@@
  • We explore how to infer the potential $w$ arising from the conjugate factorization above using only samples from $\rho$, by relying on optimal transport (OT) theory (Santambrogio, 2015). Indeed, since the condition $\rho=\nabla w^{*}\,\sharp\,e^{-w}$ is equivalent to stating that $w^{*}$ is the Brenier potential linking $e^{-w}$ to $\rho$, we can parameterize $w$ as an input-convex neural network $w_{\theta}$ whose gradient is an OT map. We provide illustrations on simple generative examples.
@@ 2 Background
We present neural approaches to solve the OT problem and introduce moment measures.
2.1 Optimal Transport
Let $\mathcal{P}(\mathbb{R}^{d})$ be the set of Borel probability measures on $\mathbb{R}^{d}$. Let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$ and set $\Pi(\mu,\nu)$ to be the set of couplings between $\mu$ and $\nu$. We consider the primal OT problem with squared-Euclidean cost,
$\mathcal{W}_{2}^{2}(\mu,\nu):=\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\tfrac{1}{2}\|x-y\|^{2}\ \mathrm{d}\pi(x,y)\,.$ (1)
This problem admits a dual formulation:
$\begin{split}&\sup\limits_{(f,g)\in\textrm{L}^{1}(\mu)\times\textrm{L}^{1}(\nu)}\int_{\mathbb{R}^{d}}f(x)\mathrm{d}\mu(x)+\int_{\mathbb{R}^{d}}g(y)\mathrm{d}\nu(y),\\ &\textrm{s.t.}\;f(x)+g(y)\leq\tfrac{1}{2}\|x-y\|^{2},\forall(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}\,.\end{split}$ (2)
The functions $f$ and $g$ that maximize the expression in equation 2 are referred to as the Kantorovich potentials. Equation 1 can itself be seen as a relaxation of the Monge formulation of the OT problem,
@@ 2.1 Optimal Transport
$\min_{\begin{subarray}{c}u\in\mathcal{C}(\mathbb{R}^{d})\\ u(0)=0\end{subarray}}\!\!\mathcal{T}_{u}(\mu,\nu):=\!\int_{\mathbb{R}^{d}}\!u(x)\,\mathrm{d}\mu(x)+\!\int_{\mathbb{R}^{d}}\!u^{*}(x)\,\mathrm{d}\nu(x)$ (4)
where $u^{*}(y):=\max_{x\in\mathbb{R}^{d}}\langle x,y\rangle-u(x)$ is the convex-conjugate of $u$ and noting that we lift any ambiguity on $u$ by selecting the unique potential such that $u(0)=0$. By denoting $\mathcal{T}(\mu,\nu)$ the optimal value and $\mathfrak{B}(\mu,\nu)$ the Brenier potential solving equation 4 above, this results in $\nu=\nabla\mathfrak{B}(\mu,\nu)\,\sharp\,\mu$.
Neural OT solvers. The goal of neural OT solvers is to estimate $\mathfrak{B}(\mu,\nu)$ using samples drawn from the source $\mu$ and the target distribution $\nu$. Makkuva et al. (2020); Korotin et al. (2020) proposed methods that leverage input convex neural networks (ICNN), originally introduced by Amos et al. (2017), to parameterize the potential $u$ as an ICNN. A key challenge in these approaches is managing the Legendre transform $u^{*}$. To overcome this, a surrogate network can be employed to approximate $u^{*}$ with recent advances by Amos (2023) improving these implementations through amortized optimization.
2.2 Moment Measures
For a given convex function $u$, the moment measure of $u$ is defined as the pushforward measure $\nabla u\,\sharp\,\mathfrak{P}_{u}$ where $\mathfrak{P}_{u}$ is the (Gibbs) log-concave probability measure with density
$\mathfrak{P}_{u}(x):=\frac{e^{-u(x)}}{\int_{\mathbb{R}^{d}}e^{-u(z)}\mathrm{d}z}\,.$ (5)
The moment measure of $u$ is in particular well defined when the quantity $\int_{\mathbb{R}^{d}}e^{-u(z)}\mathrm{d}z$ is positive and finite.
While any measure $\rho$ supported on a compact set $K$ is guaranteed to be the moment measure of some convex potential $u$ (Santambrogio, 2016), Cordero-Erausquin and Klartag showed that $u$ is often discontinuous at the boundary $\partial K$. In Cordero-Erausquin and Klartag (2015), they studied the variational problem
@@ 2.2 Moment Measures
and proved that if $\rho$ is a probability measure with a finite first moment, whose barycenter is at 00 and is not supported on a hyperplane, then it is the moment measure of an essentially continuous potential. Subsequently, Santambrogio (2016) derived the same result using a variational formulation over probability measures, given by
$\min_{\xi\in\mathcal{P}(\mathbb{R}^{d})}\{\mathcal{E}(\xi)+\mathcal{T}(\xi,\rho):\xi\in\mathcal{P}_{1}(\mathbb{R}^{d})\},$ (7)
where $\mathcal{E}$ denotes the negative Shannon entropy of a measure, and $\mathcal{T}(\xi,\rho)$ refers to the optimal value of equation 4. When $\rho$ meets the conditions enunciated in Cordero-Erausquin and Klartag (2015)Santambrogio showed that the minimizers of equation 7 coincide with the Gibbs distributions $\mathfrak{P}_{u}$, where $u$ are the convex functions whose moment measure is $\rho$.
3 Conjugate Moment Measures
We denote the convex function $u$ whose moment measure is $\rho$ as the moment potential of the distribution $\rho$. Additionally, we refer to the Gibbs distribution associated with such potential, $\mathfrak{P}_{u}$ as in equation 5, as the Gibbs factor of $\rho$.
@@ 3.1 Limitations of Moment Measures
The general case
By rewriting the KL divergence between $\rho$ and $\mathfrak{P}_{u^{*}}$ as follows:
$\textrm{KL}(\rho\|\mathfrak{P}_{u^{*}})=\mathcal{E}(\rho)+\int u^{*}\mathrm{d}\rho+\ln\left(\int e^{-u^{*}}\right),$
and applying this in the variational problem studied by Cordero-Erausquin and Klartag as referenced in equation 6, we obtain that the moment potentials of $\rho$ minimize the following problem:
@@
The quantity $\int e^{-u}\int e^{-u^{*}}$ is intrinsic to the density $\mathfrak{P}_{u}$, and has been investigated by Ball (1986) and Artstein-Avidan et al. (2004) who demonstrated that, up to a translation, this term is maximized when $\mathfrak{P}_{u}$ is Gaussian. This reformulation reveals that the Gibbs distributions associated with the Legendre transform of $\rho$’s moment potentials, $\mathfrak{P}_{u^{*}}$, are the closest log-concave distribution to $\rho$, adjusted by the regularization term $\int e^{-u}\int e^{-u^{*}}$. As a result, the Gibbs factor of $\rho$ may differ significantly from $\rho$, especially in cases where $\rho$ is a Gaussian distribution.
Gaussian distributions. The multivariate normal distribution $\mathcal{N}(m,\Sigma)$ with $\Sigma$ non-degenerate is recovered as by $\mathfrak{P}_{v}$ with $v(x)=\frac{1}{2}(x-m)^{T}\Sigma^{-1}(x-m)$. When $m\neq 0$ or $\Sigma$ is degenerate, Cordero-Erausquin and Klartag show that $\mathcal{N}(m,\Sigma)$ is not a moment measure. Yet, when $\Sigma$ is non-degenerate, moment potentials of $\mathcal{N}(0,\Sigma)$ exist and have an analytical expression given by the following proposition.
@@
Let $\rho=\mathcal{N}(0_{\mathbb{R}^{d}},\Sigma)$. If $\Sigma$ is non degenerate, the moment potential of $\rho$ is
$u_{m}(x)=\frac{1}{2}(x-m)^{T}\Sigma(x-m),$
for some $m\in\mathbb{R}^{d}$. The Gibbs factor of $\rho$ is $\mathcal{N}(m,\Sigma^{-1})$.
The proof is provided in the Appendix and relies on the fact that the OT map between two Gaussians is known in closed form Peyré et al. (2019). Intuitively this results can be interpreted as follows. When a peaked distribution $\rho$ is centered around the origin (e.g. $\Sigma\approx 0$), its corresponding moment potential is such that the image set $\nabla u$ on the support of $\mathfrak{P}_{u}$ is necessarily tightly concentrated around 00. This has the implication that $u$ is a slowly (almost constant) varying potential on the entire support of $\mathfrak{P}_{u}$. As a result, one has the (perhaps) counter-intuitive result that the more peaky $\rho$, the more spread-out $\mathfrak{P}_{u}$ must be. From this simple observation, validated experimentally in Section 3.3 we draw the intuition that a change is needed to reverse this relationship, while still retaining the interest of a measure factorization result.
@@
3.2 Adding a Conjugate
Our main theorem is stated as follows:

### v2

@@ Abstract
The canonical approach in generative modeling is to split model fitting into two blocks: define first how to sample noise (e.g. Gaussian) and choose next what to do with it (e.g. using a single map or flows). We explore in this work an alternative route that ties sampling and mapping. We find inspiration in moment measures (Cordero-Erausquin and Klartag, 2015), a result that states that for any measure $\rho$, there exists a unique convex potential $u$ such that $\rho=\nabla u\,\sharp\,e^{-u}$. While this does seem to tie effectively sampling (from log-concave distribution $e^{-u}$) and action (pushing particles through $\nabla u$), we observe on simple examples (e.g., Gaussians or 1D distributions) that this choice is ill-suited for practical tasks. We study an alternative factorization, where $\rho$ is factorized as $\nabla w^{*}\,\sharp\,e^{-w}$, where $w^{*}$ is the convex conjugate of a convex potential $w$. We call this approach conjugate moment measures, and show far more intuitive results on these examples. Because $\nabla w^{*}$ is the Monge map between the log-concave distribution $e^{-w}$ and $\rho$, we rely on optimal transport solvers to propose an algorithm to recover $w$ from samples of $\rho$, and parameterize $w$ as an input-convex neural network. We also address the common sampling scenario in which the density of $\rho$ is known only up to a normalizing constant, and propose an algorithm to learn $w$ in this setting.
@@ 1 Introduction
A decade after the introduction of GANs (Goodfellow et al., 2014) and VAEs (Kingma and Welling, 2014), the field of generative modeling has grown into one of the most important areas of research in machine learning. Both canonical approaches follow the template of learning a transformation that can map random codes to meaningful data. These transformations can be learned in supervised manner, as in the dimensionality reduction pipeline advocated in VAEs, or distributionally as advocated in the purely generative literature. In the latter, the variety of such transforms has gained remarkably in both complexity, using increasingly creative inductive biases, from Neural-ODEs(Chen et al., 2018; Grathwohl et al., 2018), diffusions (Song et al., 2020; Ho et al., 2020), optimal transport (Korotin et al., 2020) to flow-matching (Lipman et al., 2023; Tong et al., 2023; Pooladian et al., 2023).
Sample first, move next. All of these approaches are grounded, however, on choosing a standard Gaussian multivariate distribution to sample noise/codes. For both GANs and VAEs, that choice is usually made because of its simplicity, but also, in the case of diffusion models, because Gaussian distributions can be recovered quickly with suitable stochastic processes (Ornstein–Uhlenbeck). Other works have considered optimizing prior noise distributions (Tomczak and Welling, 2018; Lee et al., 2021; Liang et al., 2022) but still do so in a two step approach where mappings are estimated independently. We explore in this work a new generative paradigm where sampling and transforms are not treated separately, but seen as two facets of the same single convex potential.
Tying sampling and action. Our work looks to moment measures Cordero-Erausquin and Klartag (2015); Santambrogio (2016) to find inspiration for a factorization that ties both sampling and action.  (Cordero-Erausquin and Klartag, 2015) proved that for any sufficiently regular probability measure $\rho$, one can find an essentially-continuous, convex potential $u$ such that $\rho=\nabla u\,\sharp\,e^{-u}$: essentially, any probability distribution can be recovered by first sampling points from a log-concave distribution $e^{-u}$, and moving them with $\nabla u$ where, remarkably, the same single convex potential $u$ is used twice.
@@
  • After recalling in detail the contribution of  Cordero-Erausquin and Klartag (2015) in §2, we argue in §3.1 that the moment measure factorization may not be suitable for practical tasks. For instance, in the case of Gaussian distributions, we show that if $\rho$ is Gaussian with variance $\Sigma$, the corresponding log-concave distribution $e^{-u}$ is Gaussian with variance $\Sigma^{-1}$, amplifying beyond necessary any minor degeneracy in $\Sigma$. Similarly, for univariate $\rho$, we found that the spread of $e^{-u}$ is inversely proportional to that of $\rho$. While these results may not be surprising from a theoretical perspective, they call for a different strategy to estimate a tied sample/action factorization.
@@
  • We explore how to infer the potential $w$ from the conjugate factorization using either samples from $\rho$ or its associated energy function when the density of $\rho$ is only known up to a normalizing constant, leveraging optimal transport (OT) theory (Santambrogio, 2015). Indeed, since the condition $\rho=\nabla w^{*}\,\sharp\,e^{-w}$ is equivalent to stating that $w^{*}$ is the Brenier potential linking $e^{-w}$ to $\rho$, we can parameterize $w$ as an input-convex neural network $w_{\theta}$ whose gradient is an OT map. We provide illustrations on simple generative examples.
@@ 2 Background
Optimal Transport.
Let $\mathcal{P}(\mathbb{R}^{d})$ denote the set of Borel probability measures on $\mathbb{R}^{d}$. For $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$, let $\Pi(\mu,\nu)$ denote the set of couplings between $\mu$ and $\nu$. We consider the primal OT problem,
$\mathcal{W}_{2}^{2}(\mu,\nu):=\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\tfrac{1}{2}\|x-y\|^{2}\ \mathrm{d}\pi(x,y)\,,$ (1)
with squared-Euclidean cost. This problem admits a dual formulation:
$\begin{split}&\sup\limits_{(f,g)\in\textrm{L}^{1}(\mu)\times\textrm{L}^{1}(\nu)}\int_{\mathbb{R}^{d}}f\mathrm{d}\mu+\int_{\mathbb{R}^{d}}g\mathrm{d}\nu,\;\textrm{s.t.}\;f(x)+g(y)\leq\tfrac{1}{2}\|x-y\|^{2},\forall(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}\,.\end{split}$ (2)
The functions $f$ and $g$ that maximize the expression in (2) are referred to as the Kantorovich potentials. Equation (1) can itself be seen as a relaxation of the Monge formulation of the OT problem,
@@
$\min_{\begin{subarray}{c}u\in\mathcal{C}(\mathbb{R}^{d})\\ u(0)=0\end{subarray}}\,\int_{\mathbb{R}^{d}}\!u(x)\,\mathrm{d}\mu(x)+\!\int_{\mathbb{R}^{d}}\!u^{*}(x)\,\mathrm{d}\nu(x)$ (4)
where $u^{*}(y):=\max_{x\in\mathbb{R}^{d}}\langle x,y\rangle-u(x)$ is the convex-conjugate of $u$ and noting that we lift any ambiguity on $u$ by selecting the unique potential such that $u(0)=0$. By denoting $\mathfrak{B}(\mu,\nu)$ the Brenier potential solving problem (4) above, this results in $\nu=\nabla\mathfrak{B}(\mu,\nu)\,\sharp\,\mu$.
Neural OT solvers. The goal of neural OT solvers is to estimate $\mathfrak{B}(\mu,\nu)$ using samples drawn from the source $\mu$ and the target distribution $\nu$. Makkuva et al. (2020); Korotin et al. (2020) proposed methods that leverage input convex neural networks (ICNN), originally introduced by Amos et al. (2017), to parameterize the potential $u$ as an ICNN. A key challenge in these approaches is handling the Legendre transform $u^{*}$. To overcome this, a surrogate network can approximate $u^{*}$ with recent advances by Amos (2023) improving these implementations through amortized optimization.
Moment Measures.
For a given convex function $u$, the moment measure of $u$ is defined as the pushforward measure $\nabla u\,\sharp\,\mathfrak{P}_{u}$ where $\mathfrak{P}_{u}$ is the (Gibbs) log-concave probability measure with density $\mathfrak{P}_{u}(x):=\frac{e^{-u(x)}}{\int_{\mathbb{R}^{d}}e^{-u(z)}\mathrm{d}z}\,.$ The moment measure of $u$ is well defined when the quantity $\int_{\mathbb{R}^{d}}e^{-u(z)}\mathrm{d}z$ is positive and finite. While any measure $\rho$ supported on a compact set $K$ is guaranteed to be the moment measure of some convex potential $u$ (Santambrogio, 2016), Cordero-Erausquin and Klartag showed that $u$ is often discontinuous at the boundary $\partial K$. By studying the variational problem
@@
Cordero-Erausquin and Klartag (2015), proved that if $\rho$ is a measure with a finite first moment, a barycenter at 00 and is not supported on a hyperplane, then it is the moment measure of an essentially continuous potential (in the sense of Cordero-Erausquin and Klartag (2015)).
3 From Moment Measures to Conjugate Moment Measures
For a measure $\rho$, we call the convex function $u$ that satisfies $\nabla u\sharp\mathfrak{P}_{u}=\rho$ the moment potential of $\rho$. Additionally, we refer to the Gibbs distribution associated with such potential, $\mathfrak{P}_{u}$, as the Gibbs factor of $\rho$. We show in this section that moment measures are not well suited for generative modeling purposes, as they often produce a Gibbs factor that differs significantly from the target distribution $\rho$. This limitation is particularly evident in the case of Gaussian distributions, which motivates the introduction of an alternative representation: the conjugate moment measure.
@@ 3.1 Limitations of Moment Measures
By writing the KL divergence between $\rho$ and $\mathfrak{P}_{u^{*}}$ as $\textrm{KL}(\rho\|\mathfrak{P}_{u^{*}})=\int\ln(\rho)\mathrm{d}\rho+\int u^{*}\mathrm{d}\rho+\ln\left(\int e^{-u^{*}}\right)$ and applying this in the variational problem studied by Cordero-Erausquin and Klartag as referenced in (5), we obtain that the moment potentials of $\rho$ minimize the following problem:
@@
Equation (6) reveals that the Gibbs distributions associated with the Legendre transform of $\rho$’s moment potentials, $\mathfrak{P}_{u^{*}}$, are the log-concave distributions closest to $\rho$, up to the regularization term $-\ln\left(\int e^{-u}\int e^{-u^{*}}\right)$. This quantity has been studied by Ball (1986) and Artstein-Avidan et al. (2004), who showed that for convex functions $u$ whose corresponding Gibbs factor $\mathfrak{P}_{u}$ has its barycenter at the origin, the term $-\ln\left(\int e^{-u}\int e^{-u^{*}}\right)$ attains its minimum value of $-d\ln(2\pi)$ when $\mathfrak{P}_{u}$ is Gaussian. A direct consequence of (6) is that the Gibbs factor of $\rho$ can differ significantly from $\rho$, particularly when $\rho$ is Gaussian, as formalized in the following proposition:
@@
Let $\rho=\mathcal{N}(0_{\mathbb{R}^{d}},\Sigma)$. If $\Sigma$ is non degenerate, the moment potentials of $\rho$ are $u_{m}(x)=\frac{1}{2}(x-m)^{T}\Sigma(x-m),$ with $m\in\mathbb{R}^{d}$. The associated Gibbs factor of $\rho$ is $\mathfrak{P}_{u_{m}}=\mathcal{N}(m,\Sigma^{-1})$.
Proof.
The optimization problem in equation (6) is translation invariant: replacing $u$ with $x\mapsto u(x+a)$ for any $a\in\mathbb{R}^{d}$ leaves the objective unchanged. We can thus restrict to convex functions $u$ such that $\mathfrak{P}_{u}$ has barycenter at zero. In this setting, choosing $u(x)=\frac{1}{2}x^{\top}\Sigma x$ yields $\mathfrak{P}_{u}=\mathcal{N}(0,\Sigma^{-1})$ and $\mathfrak{P}_{u^{*}}=\mathcal{N}(0,\Sigma)$, which minimize the first term in (6), while the second term reaches its minimum value of $-d\ln(2\pi)$ (Artstein-Avidan et al., 2004). Hence, this choice minimizes the full objective, and translations of $u$ span the set of moment potentials of $\rho$. ∎
@@
Intuitively this results can be interpreted as follows. When a peaked distribution $\rho$ is centered around the origin (e.g. $\Sigma\approx 0$), its corresponding moment potential is such that the image set $\nabla u$ on the support of $\mathfrak{P}_{u}$ is necessarily tightly concentrated around 00. This has the implication that $u$ is a slowly (almost constant) varying potential on the entire support of $\mathfrak{P}_{u}$. As a result, one has the (perhaps) counter-intuitive result that the more peaky $\rho$, the more spread-out $\mathfrak{P}_{u}$ must be. From this simple observation–validated experimentally in Figure 2–we draw the intuition that a change is needed to reverse this relationship, while still retaining the interest of a measure factorization result.
3.2 The Conjugate Moment Measure factorization
Our main result, which establishes the conjugate moment measure factorization, is stated below: